# Why do explanations fail? A typology and discussion on failures in XAI

### v1

@@ Introduction
We call an explanation misleading when the failure results from the ML system being dysfunctional, i.e. when it fails to meet the very purpose it was designed for. We distinguish two situations, depending on whether the dysfunction comes from the prediction or the explanation. There is a risk that the explanation, however, could be accepted by the users without them being able to perceive this failure (”I understand — but should I?” in Fig. 1).
@@
First, it can occur that ML models output confident yet incorrect predictions. In this case we say the ML model is dysfunctional, raising the question of the relevance or potential misleadingness of generating explanations. Second, the XAI system is deemed dysfunctional, when it fails to meet the mathematical objectives it has been designed to satisfy: most explanations are generated to satisfy required properties pre-defined by the ML practitioner (e.g., optimize certain objectives). Some of them, such as gradient-based methods (e.g., (Selvaraju et al. 2017)), are directly defined as closed-form formulas. In this case, we consider that they are functional by design, and the resulting explanations faithful. Hence, the issues these methods may encounter fall into other types of failures, detailed later. On the other hand, some explanation methods rely on optimizing cost functions (see e.g., counterfactual example generation (Mothilal, Sharma, and Tan 2020)), and as such do not guarantee that the associated desiderata is satisfied. Often, there is no guarantee that a satisfying solution to these formalized problems exists, as discussed for instance by (Ye and Durrett 2022) for prompt-based explanations for Large Language Models. As a result, the explanation may be unfaithful to the model it aims at explaining.
@@
When an incorrect prediction is returned, although explanations can be useful for model calibration (Ye and Durrett 2022), they may be seen as potentially harmful, especially for users with low levels of awareness (Papenmeier, Englebienne, and Seifert 2019). An unfaithful explanation is also obviously problematic, as a lack of fidelity to the model may induce the user either to reject the system or to place unwarranted trust in it (Papenmeier, Englebienne, and Seifert 2019).Yet, despite its critical importance, this faithfulness is often overlooked by both users and ML practitioners (Kaur et al. 2020). Even when there is a will to control for faithfulness, the precise definition and evaluation of this notion remains elusive (Jacovi and Goldberg 2020), amplifying the challenge. We argue that assessing faithfulness is crucial, and that this evaluation should precede all other assessments of the explanation: firstly, contrary to other assessments, it is a purely technical task, allowing ML developers to conduct it independently of end-users. Secondly, it serves as a foundational step for identifying and addressing any other potential issue. In the rest of the paper, the ML system is therefore assumed to be functional, i.e. to provide accurate predictions and faithful explanations.
Contradictory explanations
The explanation failures that we group under the term contradictory explanations are characterized by conflicting information provided by one or several explainers, that are assumed to provide faithful information with respect to an accurate ML model to explain. More precisely, these inconsistencies occur when there is a contradiction between different pieces of information of the same explanation (competing explanations), between different explanations by the same explainer (unstable explanations), or different explanations by different explainers (incompatible explanations). One consequence of this limitation is that users may not understand this conflicting information, and wonder ”Why is it different?” (bottom part of Fig 1).
Competing explanations
We say that explanations are competing when several parts of the explanation are contradicting with one another.
@@
We identify two scenarios when this can happen. Often, explanations are composed of several components, interacting with each other in various ways: e.g., a feature attribution vector represents the contributions of each feature, word or pixel to the prediction. Yet, numerous works show that more complex effects such as interactions or correlations between features are often not taken into account  (Gosiewska and Biecek 2019; tsang2020does; mase2019explaining; Casalicchio, Molnar, and Bischl 2019; Hooker, Mentch, and Zhou 2021) A contradiction may thus appear between the semantic relationship of two notions, their actual correlation in the data used to train the model, and the explanation returned by the system. The second scenario is when XAI systems generate, rather than a single explanation, a set of explanations (e.g. of counterfactual examples (Mothilal, Sharma, and Tan 2020)) to provide richer insights to the user. In the counterfactual case, (Laugel et al. 2023) discuss the underlying notion of explanation diversity, which we connect here to the risk of getting competing explanations: diverse explanations generally aim to suggest alternative, hence sometimes contradictory, actions to the user.
@@
Whether coming from different explanations among a diverse set, or from diverse individual components within a given explanation, a contradiction between two pieces of information may be perceived as confusing by the user, potentially leading them to rejecting the explanation (Suffian et al. 2022). For instance (Bove et al. 2022) suggest competing explanations as one of the reasons for which users misunderstand the explanations returned by the SHAP approach. Furthermore, in a non-cooperative setting (Bordt et al. 2022) where the objectives of the user and the ML developer are not aligned, (Goethals, Martens, and Evgeniou 2023) discuss how this problem also opens up the risk of explanation manipulation through the selection of an explanation that is not in the best interest of the user.
@@
The proposed solutions can be grouped in two categories: the first one focuses on better XAI systems, either by adapting them, e.g. to take into account correlations (Aas, Jullum, and Løland 2021), or enriching them, e.g. proposing to calculate, in addition to the usual feature importance vectors, the contributions of feature interactions (Gosiewska and Biecek 2019). A second type of enrichment consists in exploiting expert knowledge to contextualize feature contributions: informing the user, usually through the XUI, may allow to rationalize some confusion that can be caused by competing explanations (Bove et al. 2022).
Unstable explanations
We call explanations unstable when there is an inconsistency, i.e. a contradiction, between explanations within a supposedly stable scenario: for instance, when producing local explanations for similar instances with similar outcomes, one might expect that the explanations should be similar as well (Jacovi and Goldberg 2020). Such explanation inconsistencies have been widely observed, in particular for post-hoc XAI methods (Alvarez-Melis and Jaakkola 2018a; Laugel et al. 2019). Similarly to competing explanations unstable explanations are perceived originally as a technical failure of the explainer. Yet, as we describe below, it can also originate from the user, or from a combination of both the system and the user.
@@
Unstable explanations are often considered as a technical failure of the explainer, viewed as a lack of robustness that needs to be fixed (Alvarez-Melis and Jaakkola 2018b; Slack et al. 2021; Mishra et al. 2021; Kindermans et al. 2019) However, other studies ascribe these inconsistencies to the model to be explained (Dombrowski et al. 2019; Ghorbani, Abid, and Zou 2019; Alvarez-Melis and Jaakkola 2018b): the local behavior of the latter may indeed vary abruptly, due to the complexity of the task being modeled, the complexity of the model itself or its lack of robustness. Faithful explanations then reflect these steep changes, leading to an apparent lack of stability. Moreover, this issue of instability may conflate with user perception of the similarity between instances and the explanations they expect as a result. This similarity, that depends on the user knowledge and possible biases, may differ from the similarity considered by the ML system. Pushing the expectation of explanation stability to its extreme, explanations generated for identical observations are anticipated to be identical. However, post-hoc model-agnostic methods, be they local or global (e.g.  (Ribeiro, Singh, and Guestrin 2016; Lundberg and Lee 2017; Wachter, Mittelstadt, and Russell 2018]; Altmann et al. 2010)), often rely on a stochastic data generation step (Zhou, Hooker, and Wang 2021; Visani et al. 2022) that may cause instability. This comes in addition to the Roshomon effect, i.e. that several equally good but potentially drastically different solutions can coexist and therefore be selected as explanations (Hancox-Li 2020).
@@
Considering that faithful explanations reflect the state of ML model, depending on its potential causes discussed above, a lack of stability can either be seen as an actual failure or as a desired characteristic: for instance more local explanations are expected to be less stable (Molnar 2020), with locality being a commonly expressed desideratum for explanations (Radensky et al. 2022). Thus, as for competing explanations, interpreting the lack of stability as a failure depends on the users needs, their knowledge of the explainer, and their perception of how similar the explanations should be: a user not being aware of the locality-stability tradeoff may see it as problematic but may not otherwise (Hancox-Li 2020). Still, instability may result in the user rejecting the explanation, or even the whole AI system, seeing it as a proof of unfairness by the model (Sharma, Henderson, and Ghosh 2020; Hickey, Di Stefano, and Vasileiou 2021), that may be abused by the organization providing the explanation (Goethals, Martens, and Evgeniou 2023).
@@
Various approaches have been proposed to address the unstability issue, depending on its source. To fix the stochastic unstability of model-agnostic post-hoc explainers, most contributions focus on algorithmic modifications of the random data generation step they rely on, e.g. fully replacing it with a deterministic one (Zhou, Hooker, and Wang 2021; Zafar and Khan 2019), or through a reweighting strategy (Shankaranarayana and Runje 2019). In cases when unstability originates from a misalignment between the user perception and the ML system representation, several works propose strategies to constrain the ML model during its training phase (Alvarez-Melis and Jaakkola 2018b; Dombrowski et al. 2019). On a different note, rather than mitigating the problem, several works propose to measure the explanation stability, arguing that it describes a notion of uncertainty that can be helpful for the user to better understand and use them (Gosiewska and Biecek 2019; Shankaranarayana and Runje 2019; Slack et al. 2021; Visani et al. 2022).
Incompatible explanations: the ”Disagreement Problem” in XAI
A third case where several explanations can contradict each other occurs when several explainers are used in the same setting. We call such contradictions incompatible explanations, they correspond to the prevalent (Krishna et al. 2022) and well known Disagreement Problem (see e.g., (Sundararajan and Najmi 2020; Okeson et al. 2021; Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Neely et al. 2021)).
@@
Assuming that the explanations are faithful and stable, the most high-level root cause of this issue comes from the fact that the task of providing explanations, in particular in the post-hoc setting, is essentially underdetermined (Bordt et al. 2022): as mentioned in the introduction, the concept of ”good” explanation is elusive and has been formalized in numerous different ways, taking into account different types of desiderata. For instance, because they rely on different assumptions, the explanations generated by LIME and SHAP are expected to differ, even if both are faithful (Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022). Similar discussions apply to global feature attribution methods (Okeson et al. 2021) or counterfactual explanations (Goethals, Martens, and Evgeniou 2023). Going further, some differences between explanations can also be attributed to discrepancies in the implementations of the supposedly same mathematical explanation objective (e.g. (Sundararajan and Najmi 2020) for Shapley value-based explanations). Finally, a common source of incompatibility could be attributed for the explanations generated using the same method, but different parameters. Indeed, XAI methods generally rely on hyperparameters that may not always be understood (if known at all) by the user, albeit heavily impacting the obtained explanations (Garreau and Luxburg 2020). As a result, explanations can differ on multiple bases, ranging from the top features being different to differences in order of importance or direction (Krishna et al. 2022).
@@
This disagreement between explanations is generally viewed as a problem (Sundararajan and Najmi 2020; Garreau and Luxburg 2020; Poyiadzi et al. 2021; Swamy et al. 2022; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Roy et al. 2022; Goethals, Martens, and Evgeniou 2023). Several user studies have noted it to be a source of confusion for users (Okeson et al. 2021; Krishna et al. 2022), resulting in their lowering trust in the system and therefore leading to a reject of the system as a whole or a questionable selection of the proposed explanations (e.g. based on method popularity (Kaur et al. 2020)). As for competing explanations, in a non-cooperative setting, this incompatibility between explanations may also be seen as a degree of freedom that a malicious AI practitioner may leverage to rationalize unfair decisions by choosing the explanation the most aligned to their objectives  (Slack et al. 2020; Aïvodji et al. 2019; Goethals, Martens, and Evgeniou 2023).
On the other hand, similarly to the unstable and competing cases, incompatible explanations, if faithful, can be viewed as an opportunity that may be leveraged for a better interaction with the system in a collaborative context. Indeed, explanation disagreement can be seen as a source of diversity, as noted by (Laugel et al. 2023; Goethals, Martens, and Evgeniou 2023), which is viewed positively and helps understanding, as empirically shown by (Bove et al. 2022) when combining counterfactual with global feature attributions. In other cases, disagreements between feature importance explanations are seen as an evidence of a lack of robustness on the side of the classifier, therefore used as a starting point for model auditing (Okeson et al. 2021) to improve its performance.
@@
Intuitively, circumventing the issue of incompatibility may be simply done by hiding disagreements (Roy et al. 2022). On the contrary, other works propose to emphasize them through interfaces (Bove et al. 2023), sometimes arguing, like for stability, that the level of disagreement may be used as a measure of uncertainty to validate parts of the explanations (Krishna et al. 2022). Using the same idea, other works propose to aggregate various explanation methods (Bhatt and Moura 2021; Pirie et al. 2023) to provide more robust explanations, or even propose to train new models that minimize this disagreement (Schwarzschild et al. 2023).
User-specific failures
This section discusses explanation failures that can occur when the users misinterpret the explanations provided by a ML system with no technical failures: regardless of their quality, these explanations have been shown to sometimes fail in their explanatory objective (Cheng et al. 2019; Wang et al. 2019). We propose here a typology of failures that originates from inconsistent users’ inferences, distinguishing between three categories graphically represented in Fig. 2 and discussed in turn below: mismatch failures when there is a contradiction between the ML explanations and the users’ expectations in term of format; counterintuitive failures when this contradiction concerns the explanation content; and biased inferences failures when cognitive biases inherent to each user interfere with the explanations. We believe that these user-specific failures should be known so XAI designers can understand the users’ mental model processes and support their interpretation of the provided explanations.

### v2

@@ 4.2 Misleading Explanations
We call an explanation misleading when the failure results from the ML system being dysfunctional, i.e. when it fails to meet the very purpose it was designed for. We distinguish two situations, depending on whether the dysfunction comes from the prediction or the explanation. There is a risk that the explanation, however, could be accepted by the users without their being able to perceive this failure.
@@
First, it can occur that ML models output confident yet incorrect predictions. Such ML models can be said dysfunctional, raising the question of the relevance or potential misleadingness of generating explanations. Second, the XAI system is deemed dysfunctional when it fails to meet the mathematical objectives it has been designed to satisfy. This is related to the notion of unfaithful explanations i.e. that fail to adequately account for the behavior the ML model they are associated to, see e.g.  (Jacovi and Goldberg 2020; Li et al. 2023).
Many explanations are not generated through the minimization of a cost function, but are instead defined as closed-form formulas. Examples include saliency maps in Computer Vision (Selvaraju et al. 2017), influence functions (Koh and Liang 2017), partial dependence plots (Friedman 2001; Goldstein et al. 2015), or formal explanations (Darwiche and Hirth 2020; Audemard, Koriche, and Marquis 2020; Marques-Silva 2024). Such explanations can be considered as functional by design, and faithful.
On the other hand, some explanation methods rely on optimizing cost functions, see e.g., counterfactual example generation Wachter, Mittelstadt, and Russell (2018); Mothilal, Sharma, and Tan (2020); Laugel et al. (2018a) or surrogate-based methods (Ribeiro, Singh, and Guestrin 2016), and as such do not guarantee that the associated desiderata is satisfied. Often, there is no guarantee that a satisfying solution to these problems exists, as discussed for instance by Ye and Durrett (2022) for prompt-based explanations for Large Language Models. As a result, the explanation may be unfaithful to the model it aims at explaining.
@@
When an incorrect prediction is returned, although explanations can be useful for model calibration (Ye and Durrett 2022), they may be seen as potentially harmful, especially for users with low levels of awareness (Papenmeier, Englebienne, and Seifert 2019). An unfaithful explanation is also obviously problematic, as a lack of fidelity to the model may induce the user either to reject the system (undertrust) or to place unwarranted trust in it (overtrust(Papenmeier, Englebienne, and Seifert 2019; Colin et al. 2022).
Solutions.
Despite its critical importance, this faithfulness is still often overlooked in practice, by both users and ML practitioners (Kaur et al. 2020). Even when there is a will to control for faithfulness, the precise definition and evaluation of this notion remains elusive (Jacovi and Goldberg 2020; Laugel et al. 2018b; Li et al. 2023) and often at odds with other desired criteria (Han et al. 2023; Agarwal, Tanneru, and Lakkaraju 2024) amplifying the challenge. Like previous works, we argue that assessing faithfulness is crucial, and that this evaluation should precede all other assessments of the explanation: firstly, contrary to other assessments, it is a purely technical task, allowing ML developers to conduct it independently of end-users. Secondly, it serves as a foundational step for identifying and addressing any other potential issue. In the rest of the paper, the ML system is therefore assumed to be functional, i.e. to provide accurate predictions and faithful explanations.
4.3 Competing Explanations
We propose to name explanations competing when several parts of the explanation are contradicting with one another, if for instance they consist of several counterfactual examples.
@@ 4.3 Competing Explanations
We identify two scenarios when this can happen. Often, explanations are composed of several components, interacting with each other in various ways: e.g., a feature attribution vector represents the contributions of each feature, word or pixel to the prediction. Yet, numerous works show that more complex effects such as interactions or correlations between features are often not taken into account  (Gosiewska and Biecek 2019; Tsang, Rambhatla, and Liu 2020; Mase, Owen, and Seiler 2019; Casalicchio, Molnar, and Bischl 2019; Hooker, Mentch, and Zhou 2021). A contradiction may thus appear between the semantic relationship of two notions, their actual correlation in the data used to train the model, and the explanation returned by the system. The second scenario is when XAI systems generate, rather than a single explanation, a set of explanations, e.g. of counterfactual examples (Mothilal, Sharma, and Tan 2020) to provide richer insights to the user. In the counterfactual case, we connect the underlying notion of explanation diversity (see e.g. Laugel et al. (2023)) to the risk of getting competing explanations: these diverse explanations generally aim at suggesting the user various alternatives, i.e. a choice between several possible actions, but they may appear contradictory.
@@
Contradictions between two pieces of information may be perceived as confusing by the user, potentially leading them to reject the explanation (Suffian et al. 2022). For instance Bove et al. (2022) suggest competing explanations as one of the reasons for which users misunderstand the explanations returned by SHAP. Furthermore, in a non-cooperative setting (Bordt et al. 2022) where the objectives of the user and the ML developer are not aligned, this problem can also open up the risk of explanation manipulation through the selection of an explanation that is not in the best interest of the user, see e.g. Goethals, Martens, and Evgeniou (2023); Zhou and Joachims (2023) .
@@
The proposed solutions can be grouped in two categories: the first one focuses on better XAI systems, either by adapting them to take into account correlations (Aas, Jullum, and Løland 2021; Salih et al. 2024), or enriching them, e.g. by computing, in addition to the usual feature importance vectors, feature interactions (Gosiewska and Biecek 2019; Jiang et al. 2025). A second type of enrichment consists in exploiting expert knowledge to contextualize feature contributions: informing the user, usually through the XUI, may allow to rationalize some confusion that can be caused by competing explanations (Bove et al. 2022).
4.4 Unstable Explanations
We call explanations unstable when there is an inconsistency, i.e. a contradiction, between explanations within a supposedly stable scenario: for instance, when producing local explanations for similar instances with similar outcomes, one might expect that the explanations should be similar as well (Jacovi and Goldberg 2020). Such explanation inconsistencies have been widely observed  (Alvarez-Melis and Jaakkola 2018a; Laugel et al. 2019; Yeh et al. 2019). Similarly to competing explanations unstable explanations are perceived originally as a technical failure of the explainer. Yet, as we describe below, it can also originate from the user, or from a combination of both the system and the user.
@@ 4.4 Unstable Explanations
Unstable explanations are often considered as a technical failure of the explainer, viewed as a lack of robustness that needs to be fixed (Alvarez-Melis and Jaakkola 2018b; Slack et al. 2021b; Mishra et al. 2021; Kindermans et al. 2019). However, these inconsistencies can also be ascribed to the model to be explained (Dombrowski et al. 2019; Ghorbani, Abid, and Zou 2019; Alvarez-Melis and Jaakkola 2018b): the local behavior of the latter may indeed vary abruptly, due to the complexity of the task being modeled, the complexity of the model itself or its lack of robustness. Faithful explanations then reflect these steep changes, leading to an apparent lack of stability. Moreover, this issue of instability may conflate with user perception of the similarity between instances and the explanations they expect as a result. This similarity, that depends on the user knowledge and possible biases, may differ from the similarity considered by the ML system. Pushing the expectation of explanation stability to its extreme, explanations generated for identical observations are anticipated to be identical. However, post-hoc model-agnostic methods, be they local or global (e.g.  Ribeiro, Singh, and Guestrin (2016); Lundberg and Lee (2017); Wachter, Mittelstadt, and Russell (2018); Altmann et al. (2010)), often rely on a stochastic data generation step (Zhou, Hooker, and Wang 2021; Visani et al. 2022) that may cause instability. This comes in addition to the Roshomon effect, i.e. that several equally good but potentially drastically different solutions can coexist and therefore be selected as explanations (Hancox-Li 2020).
@@
Considering that faithful explanations reflect the state of ML model, depending on its potential causes discussed above, a lack of stability can either be seen as an actual failure or as a desired characteristic: for instance more local explanations are expected to be less stable (Molnar 2020; Yeh et al. 2019), with locality being a commonly expressed desideratum for explanations (Radensky et al. 2022). Thus, as for competing explanations, interpreting the lack of stability as a failure depends on the users needs, their knowledge of the explainer and their perception of how similar the explanations should be: a user not being aware of the locality-stability trade-off may see it as problematic but may not otherwise (Hancox-Li 2020). Still, instability may result in the user rejecting the explanation, or even the whole AI system, seeing it as a proof of unfairness by the model (Sharma, Henderson, and Ghosh 2020; Hickey, Di Stefano, and Vasileiou 2021), that may be abused by the organization providing the explanation (Goethals, Martens, and Evgeniou 2023; Zhou and Joachims 2023).
@@
Various approaches have been proposed to address instability, depending on its source. To fix the stochastic instability of model-agnostic post-hoc explainers, most contributions focus on algorithmic modifications of the random data generation step they rely on, e.g. replacing it with a deterministic one (Zhou, Hooker, and Wang 2021; Zafar and Khan 2019), or through a reweighting strategy (Shankaranarayana and Runje 2019; Yeh et al. 2019). When instability originates from a misalignment between the user perception and the ML system representation, several works propose strategies to constrain the ML model during its training phase (Alvarez-Melis and Jaakkola 2018b; Dombrowski et al. 2019). On a different note, rather than mitigating the problem, several works propose to measure the explanation stability, arguing that it describes a notion of uncertainty that can be helpful for the user to better understand and use them (Gosiewska and Biecek 2019; Shankaranarayana and Runje 2019; Slack et al. 2021b; Visani et al. 2022).
4.5 Incompatible Explanations: the ”Disagreement Problem” in XAI
A third contradiction case occurs when several explainers are used in the same setting. We call them incompatible explanations, they correspond to the prevalent (Krishna et al. 2022) and well known Disagreement Problem, see e.g. Sundararajan and Najmi (2020); Han, Srinivas, and Lakkaraju (2022); Bordt et al. (2022); Neely et al. (2021).
@@ 4.5 Incompatible Explanations: the ”Disagreement Problem” in XAI
Assuming that the explanations are faithful and stable, the most high-level root cause of this issue comes from the fact that the task of providing explanations, in particular in the post-hoc setting, is essentially underdetermined (Bordt et al. 2022): as mentioned in the introduction, the concept of ”good” explanation is elusive and has been formalized in numerous different ways, taking into account different types of desiderata. For instance, because they rely on different assumptions, LIME and SHAP explanations are expected to differ, even if both are faithful (Poyiadzi et al. 2021; Han, Srinivas, and Lakkaraju 2022). Similar discussions apply to global feature attribution methods (Okeson et al. 2021) or counterfactual explanations (Goethals, Martens, and Evgeniou 2023). Going further, some differences between explanations can be attributed to discrepancies in the implementations of the supposedly same mathematical explanation objective, see e.g. Sundararajan and Najmi (2020) for the case of Shapley value-based explanations. Finally, a source of incompatibility can be attributed for the explanations generated using the same method, but different parameters. Indeed, XAI methods generally rely on hyperparameters that may not always be understood (if known at all) by the user, albeit heavily impacting the obtained explanations (Garreau and Luxburg 2020; Bansal, Agarwal, and Nguyen 2020). As a result, explanations can differ on multiple bases, for instance, in the case of feature score explanation, ranging from the top features being different to differences in order of importance or direction (Krishna et al. 2022).
@@
This disagreement between explanations is generally viewed as a problem (Sundararajan and Najmi 2020; Garreau and Luxburg 2020; Poyiadzi et al. 2021; Swamy et al. 2022; Han, Srinivas, and Lakkaraju 2022; Bordt et al. 2022; Roy et al. 2022; Goethals, Martens, and Evgeniou 2023). Several user studies have noted it to be a source of confusion for users (Okeson et al. 2021; Krishna et al. 2022), resulting in their lowering trust in the system and therefore possibly leading to a reject of the system as a whole or a questionable selection of the proposed explanations, e.g. based on method popularity (Kaur et al. 2020). As for competing explanations, in a non-cooperative setting, this incompatibility between explanations may be leveraged by a malicious AI practitioner to rationalize unfair decisions by choosing the explanation most aligned to their objectives  (Slack et al. 2020; Aïvodji et al. 2019; Goethals, Martens, and Evgeniou 2023).
On the other hand, similarly to the unstable and competing cases, incompatible explanations, if faithful, can be viewed as an opportunity that may be leveraged for a better interaction with the system in a collaborative context. Indeed, explanation disagreement can be seen as a source of diversity, as noted by Laugel et al. (2023); Goethals, Martens, and Evgeniou (2023), which is viewed positively and helps understanding, as empirically shown by Bove et al. (2022) when combining counterfactual with global feature attributions. Other works have also leveraged explanation disagreement to reduce user overreliance to the model (Reingold, Shen, and Talati 2024). In other cases, disagreements between feature importance explanations are seen as an evidence of a lack of robustness on the side of the classifier, therefore used as a starting point for model auditing (Okeson et al. 2021) to improve its performance.
@@
Intuitively, circumventing the issue of incompatibility may be simply done by hiding disagreements (Roy et al. 2022). On the contrary, other works propose to emphasize them through interfaces (Bove et al. 2023), sometimes arguing, like for stability, that the level of disagreement may be used as a measure of uncertainty to validate parts of the explanations (Krishna et al. 2022). Using the same idea, other works propose to aggregate various explanation methods (Bhatt and Moura 2021; Pirie et al. 2023; Decker et al. 2024) to provide more robust explanations, or even propose to train new models that minimize this disagreement (Schwarzschild et al. 2023).
5 Proposed Typology: User-specific
XAI Failures
This section discusses explanation failures that can occur when users misinterpret the explanations provided by a ML system with no technical failures: regardless of their quality, these explanations have been shown to sometimes fail in their explanatory objective (Cheng et al. 2019; Wang et al. 2019). We propose here a typology of failures that originates from inconsistent users’ inferences, distinguishing between three categories graphically represented in Fig. 2 and discussed in turn below: mismatch failures when there is a contradiction between the ML explanations and the users’ expectations in term of format; counterintuitive failures when this contradiction concerns the explanation content; and biased inferences failures when cognitive biases inherent to each user interfere with the explanations. We believe that these user-specific failures should be known so XAI designers can understand the users’ mental model processes and support their interpretation of the provided explanations.